# Near-Optimal Approximation of Matrix Functions by Lanczos-FAN. Amsel, T. Chen, A. Greenbaum, C. Musco, and C. Musco

### v2

@@ 1.3 Our contributions
In Section 2, we prove near instance optimality for a broad class of rational functions (Theorem 2.1). To the best of our knowledge, these are the first true instance optimality guarantees to be proven for any function besides $f(x)=1/x$. In Subsection 2.7, we discuss how results of this kind can be used to derive approximation guarantees for functions that are uniformly well approximated by rational functions. In Section 3, we prove spectrum optimality results for two especially important matrix functions: square root (Theorem 3.2) and inverse square root (Theorem 3.1). In Section 4, we present experimental evidence showing that, for many natural problem instances, our bounds are significantly sharper than the standard bound of Fact 1. We also comment on the tightness of our results with respect to their dependence on $f$ and the condition number $\kappa(\mathbf{A})$. We conclude with a discussion of next steps and open problems in Section 5.
2 Near-optimality for rational functions
@@
Since $z_{j}\not\in\mathcal{I}$, one of ±$\pm(\mathbf{A}-z_{j}\mathbf{I})$ is positive definite. For convenience, we define
@@
Our main result on rational functions is the following near instance optimality bound, which holds under a mild assumption on the number of iterations $k$:
Theorem 2.1.
Let $r(x)={n(x)}/{m(x)}$ be a degree $(p,q)$-rational function as in Eq. 8 and define $\mathbf{A}_{j}$ as in Eq. 9. Then, if $k>\max\{p,q-1\}$, the Lanczos-FA iterate satisfies the bound
@@
Above, $\kappa(\mathbf{A}_{i})$ is the condition number of $\mathbf{A}_{i}$, the ratio of the largest to smallest magnitude eigenvalues. Theorem 2.1 shows that Lanczos-FA used to approximate $r(\mathbf{A})\mathbf{b}$ satisfies Definition 1.1 with
$C=q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q}),\qquad c=1-\frac{q-1}{k}.$
To the best of our knowledge, this is the first near-instance optimality guarantee for Lanczos-FA proven for a class of functions with multiple parameters. [^4] Disregarding the trivial case when $f$ is a polynomial of degree $<k$, and thus Lanczos-FA is exact (Lemma 2.2).
Importantly, while the prefactor $q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})$ may be large, it does not depend on the number of iterations $k$. We also note that
$\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})\leq\gamma^{q},$
where
$\gamma:=\left(1+\frac{\lambda_{\textup{max}}-\lambda_{\textup{min}}}{\eta}\right),\qquad\eta:=\min_{j=1,\ldots,q}\min\{|z_{j}-\lambda_{\textup{min}}|,|z_{j}-\lambda_{\textup{max}}|\};$
i.e. $\eta$ is the minimum distance between the spectrum of $\mathbf{A}$ and the closest pole of $r(x)$. When $\mathbf{A}$ is positive-definite and the poles of $r(x)$ are non-positive, then $\gamma\leq\kappa(\mathbf{A})$.
As discussed in Section 5, we expect that Theorem 2.1 can be tightened by significantly reducing the prefactor. Nevertheless, as illustrated in Fig. 2, even in its current form, the bound improves on the standard bound of Fact 1 in many natural cases; i.e., it more tightly characterizes the observed convergence of Lanczos-FA. Finally, as discussed further in Subsection 2.7, we note that, beyond rational functions being an interesting function class in their own right, Theorem 2.1 has implications for understanding the convergence of Lanczos-FA for functions like $x^{-\alpha}$ that have much better uniform rational approximations than uniform polynomial approximations.
Figure 2: Despite its large prefactor, the bound of Theorem 2.1 qualitatively captures the convergence behavior of Lanczos-FA for rational functions. It can be tighter than the standard FOV optimality bound of Fact 1, even for a moderate number of iterations $k$. We use rational approximations to $\exp(-x/10)$ and $\log(x)$ for comparison with Fig. 1; see Section 4 for more details.
@@ 2.1 Proof Sketch
Our proof leverages the near-optimality of Lanczos-FA in computing $\mathbf{A}^{-1}\mathbf{b}$ to obtain a bound for general rational functions. For illustration, consider the simplest possible rational function for which we are unaware of any previous near-optimality bounds: $\mathbf{A}^{-2}\mathbf{b}$. The Lanczos-FA approximation for this function is $\mathbf{Q}\mathbf{T}^{-2}\mathbf{Q}^{\mathsf{T}}\mathbf{b}=\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{b}$. Using triangle inequality and submultiplicativity, we can bound the error of the approximation as
@@
$\leq\|\mathbf{A}^{-2}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{b}\|_{2}+\|\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\|_{2}\cdot\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{b}\|_{2}.$
The first term on the right hand side can be bounded because it is the error of the $\mathbf{A}$-norm optimal approximation to $\mathbf{A}^{-2}\mathbf{b}$ (indeed, $\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}$ is the $\mathbf{A}$-norm projector onto Krylov subspace). The second term, on the other hand, is proportional to the error of the $\mathbf{A}$-norm optimal approximation to $\mathbf{A}^{-1}\mathbf{b}$; i.e. of the CG iterate. We can upper bound this optimal error for approximating $\mathbf{A}^{-1}\mathbf{b}$ in terms of the optimal error for approximating $\mathbf{A}^{-2}\mathbf{b}$ by using the fact that, if $p(x)$ is a good degree $k$ polynomial approximation to $x^{-2}$, then $xp(x)$ is a good approximation to $x^{-1}$.
@@
Our proof of Theorem 2.1 generalizes the above approach. We write the Lanczos-FA error for approximating $r(\mathbf{A})\mathbf{b}$ in terms of the error of the optimal approximations to a set of simpler rational functions, and then reduce polynomial approximation of $r$ to polynomial approximation of each of these particular functions.
2.2 Notation

### v3

@@ 1.4 Our contributions
In Section 2, we prove near instance optimality for a broad class of rational functions (Theorem 4). To the best of our knowledge, this is the first true instance-optimality guarantee for Lanczos-FA to be proven for any function besides $f(x)=1/x$. In Section 2.2, we discuss how results of this kind imply related guarantees for functions that are uniformly well-approximated by rationals, which includes many functions of interest in machine learning. In Section 3, we additionally show that Lanczos-FA satisfies a weaker version of near optimality for two crucial non-rational matrix functions—the square root and inverse square root (Theorems 6 and 7). Appendix C compares the this version of optimality to Definition 2 (near instance optimality) and to that of Fact 3. In Section 4, we present experimental evidence showing that, for many natural problem instances, our bounds are significantly sharper than the standard bound of Fact 3. These experiments also demonstrate that despite its generality, Lanczos-FA often converges significantly faster than other methods in practice. We conclude with a discussion of next steps and open problems in Section 5.
2 Near optimality for rational functions
@@
Since $z_{j}\not\in\mathcal{I}$, ±$\pm(\mathbf{A}-z_{j}\mathbf{I})$ is either positive definite or negative definite. For convenience, we define
@@
Our main result on rational functions is the following near-instance-optimality bound, which holds under a mild assumption on the number of iterations $k$:
Theorem 4.
Let $r(x)={n(x)}/{m(x)}$ be a degree $(p,q)$-rational function as in Equation 4 and define $\mathbf{A}_{j}$ as in Equation 5. Then, if $k>\max\{p,q-1\}$, the Lanczos-FA iterate satisfies the bound
@@
We prove this theorem in Appendix A. Above, $\kappa(\mathbf{A}_{i})$ is the condition number of $\mathbf{A}_{i}$, the ratio of the largest to smallest magnitude eigenvalues of $\mathbf{A}_{i}$. Theorem 4 shows that Lanczos-FA used to approximate $r(\mathbf{A})\mathbf{b}$ satisfies near instance optimality, as in Definition 2, with
$C=q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q}),\qquad c=1-(q-1)/k.$
In particular, when $\mathbf{A}$ is positive definite and each of the $z_{i}$ are negative (as is the case for rational function approximations of many functions including the square root [hale_higham_trefethen_08]), then $C\leq q\kappa(\mathbf{A})^{q}$.
As discussed in Section 5, we expect that Theorem 4 can be tightened by significantly reducing the prefactor. Nevertheless, as illustrated in Figure 2, even in its current form, the bound improves on the standard bound of Fact 3 in many natural cases; i.e., it more tightly characterizes the observed convergence of Lanczos-FA. Finally, as discussed further in Section 2.2, we note that, beyond rational functions being an interesting function class in their own right, Theorem 4 has implications for understanding the convergence of Lanczos-FA for functions like the square root which are much easier to approximate with rational functions than polynomials.
Figure 2: Despite its large prefactor, the bound of Theorem 4 qualitatively captures the convergence behavior of Lanczos-FA for rational functions. It can be tighter than the standard bound of Fact 3, even for a moderate number of iterations $k$. We use rational approximations to $\exp(-x/10)$ and $\log(x)$ for comparison with Figure 1; see Section 4 for more details.
@@ 2.1 Proof Sketch
Our proof of Theorem 4 leverages the near optimality of Lanczos-FA in computing $\mathbf{A}^{-1}\mathbf{b}$ to obtain a bound for general rational functions. For illustration, consider the simplest possible rational function for which we are unaware of any previous near-optimality bounds: $\mathbf{A}^{-2}\mathbf{b}$ when $\mathbf{A}$ is positive definite. The Lanczos-FA approximation for this function is $\mathbf{Q}\mathbf{T}^{-2}\mathbf{Q}^{\mathsf{T}}\mathbf{b}=\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{b}$. Using the triangle inequality and submultiplicativity, we can bound the error of the approximation as
@@
$\leq\|\mathbf{A}^{-2}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{A}^{-2}\mathbf{b}\|_{2}+\|\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\|_{2}\cdot\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{A}^{-1}\mathbf{b}\|_{2}.$ (6)
Using the normal equations and the fact that $\mathbf{Q}$ is a basis of the Krylov subspace, it is possible to show that $\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}=\mathbf{Q}(\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q})^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}$ is the $\mathbf{A}$-norm [^2] The $\mathbf{A}$-norm is defined for positive definite $\mathbf{A}$ by $\|\mathbf{x}\|_{\mathbf{A}}=\|\mathbf{A}^{1/2}\mathbf{x}\|_{2}$. projector onto Krylov subspace. Hence
$\|\mathbf{A}^{-2}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{A}^{-2}\mathbf{b}\|_{\mathbf{A}}$ $=\min_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\mathbf{A}^{-2}\mathbf{b}-\mathbf{x}\|_{\mathbf{A}}$
$:=\min_{\deg(p)<k}\|\mathbf{A}^{-2}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{\mathbf{A}}.$
Therefore, using $\|\mathbf{x}\|_{2}\leq(1/\sqrt{\lambda_{\textup{min}}})\|\mathbf{x}\|_{\mathbf{A}}$ and $\|\mathbf{x}\|_{\mathbf{A}}\leq\sqrt{\lambda_{\textup{max}}}\|\mathbf{x}\|_{2}$, the first term on the right hand side of Equation 6 can be bounded as
$\|\mathbf{A}^{-2}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{A}^{-2}\mathbf{b}\|_{2}\leq\sqrt{\kappa(\mathbf{A})}\min_{\deg(p)<k}\|\mathbf{A}^{-2}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Similarly, the second factor of the second term on the right hand side of Equation 6 can be bounded as
$\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{A}^{-1}\mathbf{b}\|_{2}\leq\sqrt{\kappa(\mathbf{A})}\min_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Then, since $xp(x)$ is polynomial of one degree larger than $p(x)$,
$\min_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}$ $\leq\min_{\deg(p)<k-1}\|\mathbf{A}\mathbf{A}^{-2}\mathbf{b}-\mathbf{A}p(\mathbf{A})\mathbf{b}\|_{2}$
$\leq\lambda_{\textup{max}}\min_{\deg(p)<k-1}\|\mathbf{A}^{-2}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Plugging the above bounds into Equation 6 and using the fact that the eigenvalues of $\mathbf{T}$ are contained in $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ so that $\|\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\|_{2}\leq\|\mathbf{T}^{-1}\|_{2}\leq 1/\lambda_{\textup{min}}$,
$\|\mathbf{A}^{-2}\mathbf{b}-\mathsf{lan}_{k}(x^{-2};\mathbf{A},\mathbf{b})\|_{2}\leq 2\kappa(\mathbf{A})^{3/2}\min_{\deg(p)<k-1}\|\mathbf{A}^{-2}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Bounding $\kappa(\mathbf{A})^{3/2}$ by $\kappa(\mathbf{A})^{2}$ gives the bound in Theorem 4.
@@
Our proof of Theorem 4 generalizes the above approach. We write the Lanczos-FA error for approximating $r(\mathbf{A})\mathbf{b}$ in terms of the error of the optimal approximations to a set of simpler rational functions, and then reduce polynomial approximation of $r(x)$ to polynomial approximation of each of these particular functions.
2.2 Implications for non-rational functions
Theorem 4 can be used to derive guarantees for other (non-rational) functions. In particular, consider any function $f(x)$ that is uniformly well approximated by a low-degree rational function $r(x)$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$, the interval containing all of the eigenvalues of $\mathbf{A}$; i.e., suppose $\|r-f\|_{\mathcal{I}}:=\max_{x\in\mathcal{I}}|r(x)-f(x)|$ is small. A natural way to approximate $f(\mathbf{A})\mathbf{b}$, used in [aune_eidsvik_pokern_13, aune_simpson_eidsvik_13, pleiss_jankowiak_eriksson_damle_gardner_20], is to construct $r(x)$ and output $r(\mathbf{A})\mathbf{b}$, using some iterative linear solver to quickly apply the denominators of the partial fraction decomposition of $r(x)$. Alternatively, if we have an instance-optimality guarantee for Lanczos-FA on rational functions like Theorem 4, we could use Lanczos-FA to compute $r(\mathbf{A})\mathbf{b}$. The following analysis shows that simply using Lanczos-FA on $f(x)$ itself cannot be much worse.
Lemma 5.
Assume the we have the following instance-optimality guarantee for rational functions:
Here, $C_{r}$ and $c_{r}$ depend on the choice of approximant $r(x)$. Then the error of Lanczos-FA on $f(x)$ is bounded as follows:
$\|f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$
$\leq\min_{r}\Big{(}(C_{r}+2)\|\mathbf{b}\|_{2}\cdot\|f-r\|_{\mathcal{I}}+C_{r}\min_{\deg(p)<c_{r}k}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}\Big{)}.$ (8)
We prove this lemma, which follows directly from the triangle inequality, in Section A.5. Compare the form of this bound to that of Definitions 2 and 13. It is close to a near-instance-optimality guarantee, except for the first term, which requires $f(x)$ to be uniformly well-approximated by a rational function $r(x)$ on $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$. This is still much stronger than Fact 3, which requires $f(x)$ to be uniformly well-approximated by a polynomial to guarantee that Lanczos-FA provides a good approximation. There are many functions with lower degree rational approximations than polynomial approximations, even when we require the rational function $r(x)$ to have poles only in $\mathbb{R}\setminus\mathcal{I}$ (as in our Theorem 4). Such rational approximations are obtainable by the Remez algorithm [trefethen_19, Chapter 24], and for many important functions they are also known explicitly. For example, a uniform polynomial approximation to the square root on a strictly positive interval $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ requires degree $\Omega(\sqrt{\lambda_{\textup{max}}/\lambda_{\textup{min}}})$ [trefethen_19, Chapter 8]. On the other hand, a uniform rational approximation can be obtained with degree only $O(\log(\lambda_{\textup{max}}/\lambda_{\textup{min}}))$ [hale_higham_trefethen_08, trefethen_19, Chapter 25]. Likewise, a uniform polynomial approximation to $\exp(-x)$ on the interval $[0,B]$ requires degree $\Omega(\sqrt{B})$ [aggarwal_alman_22], but uniform rational approximations can be constructed with no dependence on $B$ [saff_schonhage_76]. For such functions, we expect Equation 8 to be stronger than Fact 3.
Notice also that, while in Lemma 5, we assume $f(x)$ is well-approximated by a rational function, we are not required to actually construct the approximation. Indeed, since it holds for any $r(x)$, instead of fixing a rational approximation of a certain degree, Equation 8 automatically balances $\|r-f\|_{\mathcal{I}}$, which decreases as the degree grows, with $C_{r}$, which may increase as the degree grows (see Figure 4).
3 Near Spectrum Optimality for $\mathbf{A}^{\pm 1/2}\mathbf{b}$
In the previous section, we proved that Lanczos-FA is nearly instance optimal for rational functions in the sense of Definition 2. In this section, we prove that Lanczos-FA satisfies a weaker form of near optimality for two important non-rational functions: square root and inverse square root. We term this weaker form of guarantee “near spectrum optimality”. In Appendix C, we formally define this notion and compare it to Definition 2. We first state our bound for the inverse square root.
Theorem 6.
That is, Lanczos-FA applied to the inverse square root satisfies Definition 12 (“near spectrum optimality”) with
$C=\frac{3}{\sqrt{\pi k}}\kappa(\mathbf{A}),\qquad c=\frac{1}{2}.$
We prove this theorem in Appendix D. The proof relies on comparing the error of the $k$th Lanczos iterate for $x^{-1/2}$ to that of the Lanczos iterate for $x^{-1}$. First, applying a bound from [chen_greenbaum_musco_musco_22], we use the Cauchy integral formula to upper bound the error of Lanczos-FA on $x^{\pm 1/2}$ by its error on $x^{-1}$ (Lemma 16). Second, as Equation Equation 17 shows, Lanczos-FA is nearly instance optimal for the function $x^{-1}$; that is, it outputs $p(\mathbf{A})\mathbf{b}$ where $p$ is (nearly) the degree $k$ polynomial that best approximates $x^{-1}$. Third, the best degree $k$ polynomial approximation to $x^{-1}$ must have lower error than the best degree $k/2$ approximation to $x^{-1/2}$. This is because any degree $k/2$ approximation to $x^{-1/2}$ can be squared to yield a good degree $k$ polynomial approximation to $x^{-1}$. Combining these three steps upper bounds the error of the $k$th Lanczos-FA iterate for $x^{-1/2}$ by the error of the best degree $k/2$ polynomial approximation of $x^{-1/2}$.
Nearly the same argument can be used to prove spectrum optimality of Lanczos-FA for the function $\mathbf{A}^{-1/n}\mathbf{b}$ for any $n\in\mathbb{N}$ with $C=(2^{n}-1)\cdot\kappa(\mathbf{A})/\sqrt{\pi k}$. Furthermore, using Lemma 15 of Appendix C, we can convert Theorem 6 into a near-instance-optimality guarantee at the price of strong dependence of $C$ on $\mathbf{b}$. We next state our optimality result for the matrix square root.
Theorem 7.
This bound resembles Definition 12 with
$C=\frac{3\kappa(\mathbf{A})^{2}}{k^{3/2}},\qquad c=\frac{1}{2}.$
The proof of Theorem 7 is nearly the same as that of Theorem 6, and it likewise appears in Appendix D. Ideally, if $p$ is a polynomial approximation to $x^{1/2}$, we would like to claim that $(p(x)/x)^{2}$ yields a good polynomial approximation to $x^{-1}$. However, since this function is not necessarily a polynomial, we must instead use use $\left(\frac{p(x)-p(0)}{x}\right)^{2}$, which introduces the need to include $\{0\}$ in the maximization on the right-hand side.
4 Experiments
We now present several numerical experiments to assess the quality of our instance-optimality bounds, Theorems 4 and 5. Our results show that, despite the large prefactor $C$, our bounds already supersede the standard uniform approximation bound (Fact 3) in many cases. We also compare Lanczos-FA against several recently proposed algorithms for computing matrix functions with strong theoretical guarantees. We find that, in practice, Lanczos-FA performs better than all of them. We implement Lanczos-FA in high precision arithmetic using the flamp library, which is based on mpmath [mpmath], in order to mitigate any potential impacts of finite precision arithmetic and observe the convergence behavior of the algorithm beyond the standard machine precision. [^3] Code for our experiments is available at https://github.com/NoahAmsel/lanczos-optimality/tree/neurips2024_near_optimality..
In Figure 1, we compare the performance of Lanczos to the instance-optimal KSM (which we can compute by direct methods) and against Fact 3 for various matrix functions and spectra. We use three test matrices $\mathbf{A}\in\mathbb{R}^{100\times 100}$ which all have condition number $100$. The first has a uniformly-spaced spectrum, the second has a geometrically-spaced spectrum, and the third has eigenvalues that are all uniformly-spaced in a narrow interval except for ten very small eigenvalues. We compute the bound from Fact 3 using the Remez algorithm and compute the instance-optimal approximation using least squares regression onto the Krylov basis $\mathbf{Q}$. In Figure 1, as in almost all cases we tried, Lanczos-FA performs nearly as well as the instance-optimal approximation. For instance, the error is never more than a small multiple of the optimal error in the experiments we did.
To better understand Theorem 4, we compare the bound to the true convergence curve of Lanczos-FA for various rational functions in Figure 2. We also plot Fact 3 for reference. We use the same matrices and $\mathbf{b}$ vectors as in Figure 1; results are similar if $\mathbf{b}$ is instead chosen as a uniform linear combination of $\mathbf{A}$’s eigenvectors. We choose rational functions to match the functions used for Figure 1. We construct a degree 5 rational approximation to $\exp(-x/10)$ following [saff_schonhage_76]. We construct a degree 10 approximation to $\log(x)$ using the BRASIL algorithm [clemens_21] and verify that it has real poles outside the interval of the eigenvalues. Despite Theorem 4’s exponential dependence on the degree of the rational function being applied, Figure 2 shows that it matches the shape of the convergence curve well and is tighter than Fact 3 when the number of iterations is large. That said, in all cases plotted, Lanczos-FA always returns an approximation much closer to optimal than predicted by Theorem 4, suggesting that the leading coefficient in our bound is pessimistic.
4.1 Dependence on the rational function degree
Theorem 4 upper bounds the optimality ratio by $C=q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})$. We conjecture that this bound is far from tight. However, the following experiment provides evidence that it is not possible to entirely eliminate the dependence on the rational function’s denominator degree $q$. In particular, for parameters $(\kappa,q)$, consider approximating $\mathbf{A}^{-q}\mathbf{b}$ where $\mathbf{A}$ has spectrum $\lambda_{1}=1$ and $\lambda_{2},\ldots,\lambda_{100}$ evenly spaced between $0.999995\cdot\kappa$ and $\kappa$. In this case, $\kappa(\mathbf{A}_{1})=\cdots=\kappa(\mathbf{A}_{q})=\kappa(\mathbf{A})=\kappa$. We generate a grid of problems by picking different combinations of $(\kappa,q)$ and tuning $\mathbf{b}$ in a limited way to maximize the maximum ratio between the error of Lanczos-FA and the optimal Krylov error over all iterations. In particular, we took $\mathbf{b}$ to be an all ones vector, except we varied its first entry, using grid search to maximize the optimality ratio. The results, plotted in Figure 3, suggest that the optimality ratio grows at least as $\Omega(\sqrt{q\cdot\kappa(\mathbf{A})})$. We have yet to find harder problem instances than this.
Figure 3: The maximum observed ratio between the error of Lanczos-FA and the optimal error over choices of $\mathbf{b}$ when approximating $\mathbf{A}^{-q}$ for matrices with varying condition number $\kappa$. Each point corresponds to a pair $(\kappa,q)$. Points with the same color have the same value of $\kappa$. On the left, the dotted line plots $\sqrt{q\kappa}$ for the maximum $\kappa$ considered ($10^{6}$). On the right, the dotted line plots $\sqrt{q\kappa}$ for the maximum $q$ considered ($2^{6}$). Overall, the optimality ratio appears to scale at least as $\Omega(\sqrt{q\kappa})$.
4.2 Non-rational functions
As noted in Section 2.2, an instance-optimality guarantee for rational functions also implies that Lanczos-FA performs well on functions that are well-approximated by rationals. As an example, we consider the function $f(x)=x^{-0.4}$, for which a rational approximation in any degree can be found using the BRASIL algorithm [clemens_21]. Figure 4 shows how applying Lanczos-FA to these rational approximations compares to applying Lanczos-FA directly to the $f(x)$ itself to approximate $\mathbf{A}^{-0.4}\mathbf{b}$. When the number of iterations is small, both methods perform nearly optimally, as the accuracy is limited more by the small size of the Krylov subspace than by the difference between $f(x)$ and the rational approximant (that is, the second term in Equation 8 dominates the first term). As the number of iterations grows, the error due to approximating $f(x)$ in the Krylov subspace continues to decrease while the error of uniformly approximating $f(x)$ by some fixed rational function remains fixed (that is, the first term in Equation 8 dominates the first term); however, increasing the degree of the rational approximant decreases this source of error. This shows that understanding the convergence of Lanczos-FA for the entire family of rational approximations goes a long way toward explaining its convergence for non-rational functions. In the limit, Lanczos-FA applied to $f(x)$ itself appears to automatically inherit the instance optimality of a suitably high-degree rational approximation.
Figure 4: Applying Lanczos-FA to the function $\mathbf{A}^{-0.4}$ and rational approximations of various degrees found using the BRASIL algorithm [clemens_21]. In this experiment, the spectrum of $\mathbf{A}$ contains two clusters: 10 eigenvalues uniformly spaced near 1, and 90 eigenvalues uniformly spaced near 100. As predicted by the bound in Section 2.2, convergence of Lanczos-FA for this function appears to closely track that of a high degree rational approximant.
This result has an additional implication. A number of papers use explicit rational approximations to compute non-rational matrix functions [aune_eidsvik_pokern_13, aune_simpson_eidsvik_13, guttel_schweitzer_21, pleiss_jankowiak_eriksson_damle_gardner_20]. These approximations are often applied by applying conjugate gradient (or a related method) to each of the terms in the sum [guttel_schweitzer_21, pleiss_jankowiak_eriksson_damle_gardner_20]. In the case conjugate gradient is used, the resulting algorithm is mathematically identical to Lanczos-FA used to compute the the rational approximation. However, Figure 4 suggests that simply using Lanczos-FA on the original function is both simpler and converges faster (though memory usage and other factors may need to be considered).
Another line of work uses specialized iterative methods that exploit problem structure to apply the rational approximations [frostig_musco_musco_sidford_16, allen_zhu_li_17, musco2019spectrum, jin_sidford_19]. In Figure 5, we compare Lanczos-FA to two such methods from [jin_sidford_19] for computing $\operatorname{sign}(\mathbf{A})$, for $\mathbf{A}$ of the form $\mathbf{A}=\mathbf{B}^{\mathsf{T}}\mathbf{B}-\lambda\mathbf{I}$. While they achieve better theoretical bounds than are known for Lanczos-FA, Lanczos-FA far outperforms them on the test problems used in [jin_sidford_19].
Figure 5: A comparison of Lanczos-FA with two methods from [jin_sidford_19] (“rational” and “slanczos”) for computing the matrix sign function, which work by using a stochastic iterative method to approximate rational approximations to the step function of various degrees. The “rational” method is the main one studied in [jin_sidford_19], while “slanczos” is included because it is the best performing in their experiments. Each panel corresponds to one of the test problems from [jin_sidford_19]. Iterations of these methods are counted in number of inner products with rows of $\mathbf{A}$ rather than number of matrix-vector products with $\mathbf{A}$ as a whole. To compare these with Lanczos-FA, we consider $d$ such inner products to be equivalent to one matrix-vector product.
Additional Experiments.
In Section 3 we introduce bounds for the matrix square root and inverse square root, and in Section E.1 we provide numerical tests to study the sharpness of the bounds, verifying that they can improve on Fact 3. In Section E.2, we demonstrate the convergence behavior of Lanczos-FA on rational functions with poles inside the range of eigenvalues (Figure 6). This illustrates why a bound like Theorem 4 is not possible, but suggests a weaker bound, such as the bound in [chen_meurant_24] for $f(x)=1/x$ and indefinite $\mathbf{A}$, may be possible. Section E.3 shows that, unlike Lanczos-FA, a related algorithm called Lanczos-OR [chen_greenbaum_musco_musco_23] (which is exactly optimal for rational functions, though not in the Euclidean norm) can perform poorly on high degree rational functions when the error is measured in the Euclidean norm.
5 Outlook
This paper provides instance-optimality guarantees for Lanczos-FA applied to a range of rational functions. We conclude with open questions that we believe are worthy of further study.
Extension to other function classes.
Empirically, Lanczos-FA seems to be nearly instance optimal for a wide variety of functions beyond those considered in this paper, such as rational functions with conjugate pairs of complex poles whose real parts lie outside $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$. As seen in Figure 6, the error of Lanczos-FA on functions with real poles in $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ is intriguing, oscillating between being very large and nearly optimal. We discuss this more in Section E.2. It would be valuable to provide bounds explaining these behaviors.
It would be also natural to try to extend our bounds to Stieltjes/Markov functions, which can be viewed as a certain type of infinite degree rational function approximations with poles in $(-\infty,0]$, and includes important functions like the inverse square root and a shifted logartihm. Our bound Theorem 4 cannot be directly applied to this class due to the dependence on the rational function denominator degree $q$.
Figure 6: Convergence of Lanczos-FA for rational functions with poles in $\mathbf{A}$’s eigenvalue range or that are imaginary. The optimality ratio can be very large for some iterations. Similar behavior is seen for functions like $\operatorname{sign}(\mathbf{A})$ that have a discontinuity in the interval of $\mathbf{A}$’s eigenvalues. However, the “overall” convergence of Lanczos-FA still appears to closely track the instance-optimal solution.
Construction of hard instances / refined upper bounds.
Theorem 4 has an exponential dependence on the degree of the rational function’s denominator $q$, which limits the practicality of our bounds. It is unclear if and when this dependence can be improved. The experiment in Section 4.1 provides strong evidence that some dependence on $q$ is necessary, but the hardest examples we have depend on $\sqrt{q}$, instead of the current bound of $\kappa(\mathbf{A})^{q}$ guaranteed by Theorem 4. It is an open question whether Theorem 4 can be tightened, or whether matching hard instances exist.
Finite precision arithmetic.
Our analysis concerns the behavior of the Lanczos algorithm when run in exact arithmetic. In practice, the implementation of the Lanczos algorithm is very important; for instance, practical implementations often output a $\mathbf{Q}$ which is far from orthogonal [meurant_strakos_06, carson_liesen_strakos_24]. While this instability can be mitigated with more expensive implementations, theoretical work shows that, surprisingly, Lanczos and Lanczos-FA can work well despite it [paige_71, paige_76, paige_80, greenbaum_89, druskin_knizhnerman_91, druskin_greenbaum_knizhnerman_98]. For example, [druskin_knizhnerman_91, musco_musco_sidford_18, chen_24] show that Fact 3 still holds up to a close approximation in finite precision arithmetic for any bounded matrix function. It would be valuable to study whether stronger near-optimality guarantees like those proven in Theorem 4 are also robust to finite precision. For $f(x)=1/x$, this problem has been studied in [greenbaum_89].
Acknowledgments:
This research was supported in part by NSF Awards 2045590 (Chen, Ch. Musco), 2046235 (Ca. Musco), 2427363 (Chen, Ca. Musco, Ch. Musco), and 2234660 (Amsel).
[…]
Appendix A Proof of Theorem 4
A.1 Notation